OpenStudy (anonymous):

could someone please help me to solve 15x^2-7x-2, and show me all of the steps? I woiuld really appreciate the help with my last question

OpenStudy (saifoo.khan):

w(-4w^2 - 16w + 20) 4w(-w^2 - 4w + 5) -4w(w^2 + 4w - 5)

OpenStudy (anonymous):

-4w^3-16w^2+20w =-4w ( w^2 +4w -5) =-4w(w+5)(w-1)

OpenStudy (anonymous):

-a^3-100a = -a(a^2 +100)